OpenStudy (anonymous):

There is a linear relationship between the number of chirps made by the striped ground cricket and the air temperature. A least-squares fit on some data collected by a biologist gives the model y(hat) = 25.2 = 3.3x where x is the number of chirps per minute and y(hat) is the estimated temperature in degrees Fahrenheit. What is the estimated increase in temperature that corresponds to an increase of 5 chirps per minute?

OpenStudy (anonymous):

@amistre64 @Hero @agent0smith

hero (hero):

Your linear equation has two equal signs in it. So not able to tell if the second equation should have a plus or minus. Probably a plus.

hero (hero):

y = 25.2 + 3.3x Maybe that's what you meant

OpenStudy (anonymous):

sorry the equation I meant to put is y=25.5 + 3.3x

hero (hero):

Right and since: x = number of chirps her minute and x = 5 in this case, then y = 25.2 + 3.3(5) Compute the value of y.

OpenStudy (anonymous):

so y= 25.5 + 3.3(5) = 42?

hero (hero):

\(y = 25.5 + 3.3(5)\) \( = 41.7\) \(\approx 42\) Thus the increase in temperature in degrees Fahrenheit given 5 chirps per minute is estimated to be approximately \(42^{\circ}F\)